# From Flat to Hierarchical : Extracting Sparse Representations with Matching Pursuit

### v1

@@ 4.2 Representations from Pretrained Models
We analyze representations from pretrained vision-language models in the following experiments. This helps us avoid the challenge of flexibility of model representations in solely language-driven domains [73] (though see Sec. B.5 for preliminary results). Our evaluation is organized in four parts. We begin by (i) assessing MP-SAE’s expressivity relative to existing SAEs. We then (ii) analyze the structure of its learned representations through effective rank and coherence metrics. Next, (iii) we test its robustness to inference-time sparsity variation. Finally, (iv) we investigate its ability to uncover multimodal features in vision-language models that remain inaccessible to existing SAEs.
@@
Intuitively, $\mu_{1}(r)$ reflects how well a single concept can be approximated by a group of $r$ others; lower values indicate better separability. Fig. 7 reports $\mu_{1}(r)$ for both the full dictionary and for the subsets of concepts co-activated at inference time. Interestingly, MP-SAE learns dictionaries with higher overall Babel scores than standard SAEs, indicating greater global interference. However, the concepts it selects during inference exhibit lower Babel scores, reflecting MP-SAE’s tendency to construct conditionally orthogonal representations at inference (even when the full dictionary is correlated). By contrast, linear SAEs enforce global quasi-orthogonality in the dictionary but do not control which features co-activate at inference. As a result, inference often selects interfering directions despite a well-structured dictionary.
Figure 8: Reconstruction error as a function of inference-time sparsity $k$. When increasing SAEs’ sparsity at inference on DINOv2 representations, we see MP-SAE exhibits monotonic improvement, while TopK SAEs may degrade due to sparsity mismatch. Similar results emerge in other settings (see App. B.4B.5).
@@
Figure 9: MP-SAE identifies shared structures across modalities. Top: UMAP of sparse codes (yellow denotes image representations; blue denotes text). Bottom: Distribution of modality scores. We see existing SAEs yield features skewed towards a specific modality, suggesting bimodal structure in model representations. However, MP-SAE recovers multimodal units with shared activation across text and vision inputs.
@@
We evaluate whether MP-SAE can recover shared concepts from vision-language models (VLMs). Our analysis focuses on representations from CLIP [75], with similar results observed for AIMv2 [100], SigLIP [97], and SigLIP2[101]. All models are evaluated on the full COCO dataset [102], using both image and caption representations, balanced to ensure equal sampling across modalities. When trained on these joint embedding spaces, classical SAEs frequently learn split dictionaries [103, 104], where distinct features respond exclusively to either visual or textual inputs. This occurs despite the alignment of the underlying representation space and reflects a structural limitation in how these models extract shared features. To quantify modality selectivity, we use the Modality Score from [105], defined for concept $i$ as:
@@
where $\iota$ and $\tau$ denote activations over image and text inputs, respectively. Scores near 1 indicate image specificity; near 0, text specificity; and intermediate values reflect balanced, multimodal activation. Consistent with prior work [105], we find that standard SAEs produce sharply bimodal modality score distributions. In contrast, MP-SAE yields a significantly flatter distribution with substantial mass in the midrange, suggesting that it recovers genuinely multimodal units responsive to both modalities (see Fig. 9). We hypothesize that MP-SAE’s residual based inference plays a central role: once modality-specific information is explained, the residual shifts toward shared structure, allowing subsequent steps to capture cross-modal features. Prior work [104] has attempted to bridge the modality gap by applying a learned translation from one modality to the other. This can be seen as equivalent to a single inference step in MP-SAE. In contrast, MP-SAE continues this process iteratively, progressively refining the residual and revealing joint semantic structure.
@@ 5 Conclusion
In this work, we revisit SAEs to understand their ability to identify concepts encoded in a manner outside the scope of their motivating hypothesis, i.e., LRH. Our results show that while standard SAEs are indeed effective under LRH, they struggle to capture more complex representational structures—such as hierarchical, nonlinear, and multimodal features—that have emerged in recent studies of large neural networks. To contextualize these results with respect to an SAE that is motivated to (partially) accommodate such phenomenology outside LRH’s scope, we introduced MP-SAE. Specifically, an extension of the classical Matching Pursuit algorithm, MP-SAE promotes conditional orthogonality through residual-guided inference. Our results show that this design enables MP-SAE to recover rich, hierarchically organized, and nonlinearly accessible features that standard SAEs missed. Beyond interpretability, we also find that residual-based, sequential encoders like MP-SAE offer practical advantages, such as adaptive inference-time sparsity and progressive feature discovery, which may be of independent interest to the community. Overall, we argue our results highlight the limitations of assuming purely linear structure in neural representations.
@@
We emphasize that our goal with proposing MP-SAE was not to propose a “superior” SAE architecture, but to use it as a tool whose inductive bias aligns with a particular class of concept structures, namely, hierarchically and nonlinearly organized features. Under this hypothesis, MP-SAE proves more expressive, more modular, and more robust than standard SAEs. However, we note MP-SAE assumes that representations can be incrementally explained by conditionally orthogonal features. This inductive bias may be poorly matched in flat or entangled regimes, or in settings with categorical structure. Matching Pursuit is also a greedy algorithm, which lacks global optimality and may be brittle under extreme noise (though we do find existing SAEs fail in this regime as well).
@@ Acknowledgements
Authors thank the CRISP group at Harvard SEAS for insightful conversations, and the Kempner Institute and CBS-NTT program in Physics of Intelligence at Harvard University for access to compute resources used for performing experiments reported in this paper. Valérie Costa thanks the Bertarelli Foundation for supporting her work as a fellow and providing funding as part of the Bertarelli Research Award.
@@ Appendix A Related Work
Sparse coding
Sparse linear generative models aim to find a sparse representations $\bm{z}\in\mathbb{R}^{p}$ that explains the data $\bm{x}\in\mathbb{R}^{m}$ via a collection of atoms, forming an overcomplete dictionary $\bm{D}\in\mathbb{R}^{m\times p}$ ($p\gg m$.) [106, 32]. Sparse representations are ubiquitous in science and engineering, with origins from computational neuroscience [106, 32] and applications in medical imaging [107, 108, 109], image restorations [110, 111, 112], radar sensing [113], transcriptomics  [114, 115], and genomics [116, 117].
@@
Unrolling for sparse coding
To address the slow convergence of sparse coding, in 2010, learned ISTA (LISTA) [134] sparked a new approach, which is now known as unrolling, to turn and relax the iterations of ISTA into layers of a neural network to accelerate the inner problem. Since then, numerous works have theoretically studied the linear convergence of unrolled ISTA [135, 136], and some others have highlighted the implicit acceleration that this approach may bring in learning the dictionary [137, 138]. Moreover, a parallel literature has theoretically shown that approximate sparse coding, such as a shallow ReLU network [139, 140], or shallow hard-thresholding (recently named as JumpReLU [18][141, 140] is practically enough to perform dictionary learning.
Interpretability of Neural Networks
Recently,  [51] has shown that many interpretability methods used to extract concepts were in fact sparse coding methods. The field has re-emerged as a compelling framework for interpreting the internal mechanisms of high-performing, large-scale models [142, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10], particularly in light of persistent challenges surrounding safety and interpretability [11, 12, 13]. Within this renewed interest, Sparse Autoencoders (SAEs) have gained prominence as a principled method for exposing the latent structure of neural representations [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Empirical support for LRH spans both vision and language domains. In vision models, internal representations have been shown to encode factors such as geometry, lighting, facial structure, and scene composition [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 50]. In language models, analogous patterns emerge, with latent directions capturing syntactic and semantic content [52, 53, 54, 55, 56, 57], character roles, arithmetic concepts [58, 59, 60], relational structures [63, 64, 65, 66], and even safety-related behaviors such as refusal to generate harmful outputs [61, 62].
@@ Appendix B Extended Experimental Details and Further Results
B.1 Synthetic experiments
The synthetic experiments were conducted using the codebase provided by Bussmann et al. [20], which we extended to accommodate both MP-SAE and BatchTopK variants. To contextualize our experimental setup, we begin by introducing the concept of Matryoshka Sparse Autoencoders (SAEs), before detailing the design and implementation of our synthetic experiments.
@@ B.1 Synthetic experiments
B.1.1 Experiment details
Dataset
The tree structure contains 20 nodes, each corresponding to a unit-norm vector in $\mathbb{R}^{20}$, as illustrated in Figure 2 and detailed in Figure 10. The root node (index 0) is fixed to the zero vector, i.e., $0\in\mathbb{R}^{20}$. All child nodes are mutually exclusive, leading to input sparsity levels of either 1 or 2. Leaf nodes—green parents without children (indices 13 to 20) are independently activated with probability 0.05. Internal green parent nodes (1, 5, and 9) are activated with probability 0.2. Blue child nodes (2–4, 6–8, 10–12) are activated with probability 0.2, conditional on their corresponding parent (1, 5, or 9) being active. This results in an expected target sparsity of $\ell_{0}=1.36$. If a node is active, its sparse code is drawn from a Gaussian distribution $z\sim\mathcal{N}(1.5,0.25^{2})$, and the input $\bm{x}$ is constructed as the sum of the active node vectors scaled by their corresponding code values. The exact tree configuration used in our experiments is available on the project’s GitHub repository222 https://github.com/mpsae/MP-SAE. Initially, all dictionary directions are orthogonal, obtained via QR decomposition. To introduce correlations specifically between nodes at the same level, we perturb each dictionary element as
@@ B.1.1 Experiment details
where the sum is restricted to other elements with the same parent, followed by renormalization. The parameter $\varepsilon$ is tuned empirically to achieve the desired level of intra-level correlation.
Training
Samples are drawn from the tree structure in batches of 200 over 15,000 training steps, using the Adam optimizer with $\beta=(0.5,0.9375)$ and a learning rate of $3\times 10^{-2}$. To stabilize optimization, gradient norms are clipped at 1.
Vanilla SAE and Matryoshka SAE are trained following the procedure described in Bussmann et al. [20]. For sparsity control, all SAEs, except MP-SAE, are trained to match a target $\ell_{0}$ sparsity of 1.36. Vanilla and Matryoshka SAEs begin with a low $\ell_{1}$ regularization during the first 3,000 steps, which is subsequently increased or decreased based on whether the observed sparsity is below or above the target.
@@
MP-SAE, in contrast, does not rely on a fixed sparsity target. Instead, it unrolls the encoder until either the residual norm drops below 0.05 or the support of the sparse code $z$ stabilizes—avoiding infinite loops when the residual cannot be sufficiently reduced. Unlike the other SAEs, MP-SAE uses tied weights.
B.1.2 Detailed Results on Synthetic Data
@@
All models are evaluated using a fixed sparse code $\bm{z}$, shown in Figure 11(b), which depicts the ground truth activation matrix $\bm{z}\in\mathbb{R}^{N\times p}$. Sparsity is maintained between 1 and 2 active units per sample. Blue child nodes (indices 2–4, 6–8, 10–12) are only activated when their corresponding green parent node (1, 5, and 9, respectively) is active, and children under the same parent are mutually exclusive.
To better understand how each model handles structured representations, we evaluate their behavior under varying levels of intra-level correlation. Vanilla SAE and BatchTopK consistently exhibit feature absorption, where child features align with those of their parents. Matryoshka SAE sometimes alleviates this issue but does so by suppressing intra-level correlations entirely, failing to capture meaningful dependencies within each level of the hierarchical tree. In contrast, MP-SAE reliably recovers both the hierarchical structure and intra-level correlations at low to moderate levels (0, 0.3). As correlation increases (0.6), MP-SAE attempts to balance both objectives, but at high correlation levels (0.9), it prioritizes preserving the hierarchy over matching correlation. We summarize our detailed observations for each correlation level below:
@@
  • Figure 15 (Correlation = 0.9): At high levels of intra-level correlation, all SAEs exhibit degraded performance. Interestingly, Vanilla SAE and BatchTopK still manage to capture some similarity among the last-layer child nodes. MP-SAE prioritizes hierarchical consistency over correlation. Rather than activating all correlated children, it selects a single child per parent, thereby sacrificing correlation to preserve the underlying conditional structure of the hierarchy.

### v2

@@ 4.2 Representations from Pretrained Models
We analyze representations from pretrained vision-language models in the following experiments. This helps us avoid the challenge of flexibility of model representations in solely language-driven domains [73] (though see Sec. B.5 for preliminary results) Our evaluation is organized in four parts. We begin by (i) assessing MP-SAE’s expressivity relative to existing SAEs. We then (ii) analyze the structure of its learned representations through effective rank and coherence metrics. Next, (iii) we test its robustness to inference-time sparsity variation. Finally, (iv) we investigate its ability to uncover multimodal features in vision-language models that remain inaccessible to existing SAEs.
@@
Intuitively, $\mu_{1}(r)$ reflects how well a single concept can be approximated by a group of $r$ others; lower values indicate better separability. Fig. 7 reports $\mu_{1}(r)$ for the learned dictionary, as well as the average over multiple representations for subsets of co-selected concepts at inference time. Interestingly, MP-SAE learns dictionaries with higher overall Babel scores than standard SAEs, indicating greater global interference. However, the concepts it selects during inference exhibit lower Babel scores, reflecting MP-SAE’s tendency to construct conditionally orthogonal representations at inference (even when the full dictionary is correlated). By contrast, linear SAEs enforce global quasi-orthogonality in the dictionary but do not control which features co-activate at inference. As a result, inference often selects interfering directions despite a well-structured dictionary.
Figure 8: Reconstruction error as a function of inference-time sparsity $k$. When increasing SAEs’ sparsity at inference on DINOv2 representations, we see MP-SAE exhibits monotonic improvement, while TopK SAEs may degrade due to sparsity mismatch. Similar results emerge in other settings (see Appendix B.5).
@@
We evaluate whether MP-SAE can recover shared concepts from vision-language models (VLMs). Our analysis focuses on representations from CLIP [75], with consistent results observed for AIMv2 [100], SigLIP [97], and SigLIP2[101] (See Appendix B.4). All models are evaluated on the COCO dataset [102], using both image and caption embeddings, balanced to ensure equal sampling across modalities. When trained on these joint embedding spaces, classical SAEs frequently learn split dictionaries [103, 104], where distinct features respond exclusively to either visual or textual inputs. This occurs despite the alignment of the underlying representation space and reflects a structural limitation in how these existing SAEs extract shared features. To quantify modality selectivity, we use the Modality Score from [105], defined for concept $i$ as:
@@
where $\iota$ and $\tau$ denote activations over image and text inputs, respectively. Scores near 1 indicate image specificity; near 0, text specificity; and intermediate values reflect balanced, multimodal activation.
Figure 9: MP-SAE identifies shared structures across modalities. Top: UMAP of sparse codes (yellow denotes image representations; blue denotes text). Bottom: Distribution of modality scores. We see existing SAEs yield features skewed towards a specific modality, suggesting bimodal structure in model representations. However, MP-SAE uniquely recovers multimodal units with shared activation across text and vision inputs.
Consistent with prior work [105], we find that standard SAEs yield sharply bimodal modality score distributions, confirming their tendency to separate modalities. In contrast, MP-SAE yields a significantly flatter distribution with substantial mass in the midrange (see Fig. 9), suggesting that it recovers genuinely multimodal units responsive to both modalities. This ability to extract multimodal concepts appears unique to MP-SAE. We hypothesize that MP-SAE’s residual based inference plays a central role: once modality-specific information is explained, the residual shifts toward shared structure, allowing subsequent steps to capture cross-modal features (See Fig. 18). Prior work [104] has attempted to bridge the modality gap by applying a learned translation from one modality to the other. This can be seen as equivalent to a single inference step in MP-SAE. However, MP-SAE continues this process iteratively, progressively refining the residual and revealing joint semantic structure. This iterative mechanism enables MP-SAE to extract hierarchical and non-linearly accessible multimodal features beyond the scope of conventional SAEs.
@@ 5 Conclusion
In this work, we revisit SAEs to understand their ability to identify concepts encoded in a manner outside the scope of their motivating hypothesis, i.e., LRH. Our results show that while standard SAEs are indeed effective under LRH, they struggle to capture more complex representational structures—such as hierarchical, nonlinearly accessible, and multimodal features—that have emerged in recent studies of large neural networks. To contextualize these results with respect to an SAE that is motivated to (partially) accommodate such phenomenology outside LRH’s scope, we introduced MP-SAE. Specifically, an extension of the classical Matching Pursuit algorithm, MP-SAE promotes conditional orthogonality through residual-guided inference. Our results show that this design enables MP-SAE to recover rich, hierarchically organized, and nonlinearly accessible features that standard SAEs missed. Beyond interpretability, we also find that residual-based, sequential encoders like MP-SAE offer practical advantages, such as adaptive inference-time sparsity and progressive feature discovery, which may be of independent interest to the community. Overall, we argue our results highlight the limitations of assuming purely linear structure in neural representations.
@@
We emphasize that our goal with proposing MP-SAE was not to propose a “superior” SAE architecture, but to use it as a tool whose inductive bias aligns with a particular class of concept structures, namely, hierarchically and nonlinearly accessible features. Under this hypothesis, MP-SAE proves more expressive, more modular, and more robust than standard SAEs. However, we note MP-SAE assumes that representations can be incrementally explained by conditionally orthogonal features. This inductive bias may be poorly matched in flat or entangled regimes, or in settings with categorical structure. Matching Pursuit is also a greedy algorithm, which lacks global optimality and may be brittle under extreme noise (though we do find existing SAEs fail in this regime as well).
@@ Acknowledgements
Authors thank the CRISP Group at Harvard SEAS for insightful conversations, and the Kempner Institute and CBS-NTT program in Physics of Intelligence at Harvard University for access to compute resources used for performing experiments reported in this paper. Valérie Costa thanks the Bertarelli Foundation for supporting her work as a fellow.
@@ Appendix A Related Work
Sparse coding Sparse linear generative models aim to find a sparse representations $\bm{z}\in\mathbb{R}^{p}$ that explains the data $\bm{x}\in\mathbb{R}^{m}$ via a collection of atoms, forming an overcomplete dictionary $\bm{D}\in\mathbb{R}^{m\times p}$ ($p\gg m$.) [106, 32]. Sparse representations are ubiquitous in science and engineering, with origins from computational neuroscience [106, 32] and applications in medical imaging [107, 108, 109], image restorations [110, 111, 112], radar sensing [113], transcriptomics  [114, 115], and genomics [116, 117].
@@
Unrolling for sparse coding To address the slow convergence of sparse coding, in 2010, learned ISTA (LISTA) [134] sparked a new approach, which is now known as unrolling, to turn and relax the iterations of ISTA into layers of a neural network to accelerate the inner problem. Since then, numerous works have theoretically studied the linear convergence of unrolled ISTA [135, 136], and some others have highlighted the implicit acceleration that this approach may bring in learning the dictionary [137, 138]. Moreover, a parallel literature has theoretically shown that approximate sparse coding, such as a shallow ReLU network [139, 140], or shallow hard-thresholding (recently named as JumpReLU [18][141, 140] is practically enough to perform dictionary learning.
Interpretability of Neural Networks Recently,  [51] has shown that many interpretability methods used to extract concepts were in fact sparse coding methods. The field has re-emerged as a compelling framework for interpreting the internal mechanisms of high-performing, large-scale models [142, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10], particularly in light of persistent challenges surrounding safety and interpretability [11, 12, 13]. Within this renewed interest, Sparse Autoencoders (SAEs) have gained prominence as a principled method for exposing the latent structure of neural representations [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Empirical support for LRH spans both vision and language domains. In vision models, internal representations have been shown to encode factors such as geometry, lighting, facial structure, and scene composition [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. In language models, analogous patterns emerge, with latent directions capturing syntactic and semantic content [52, 53, 54, 55, 56, 57], character roles, arithmetic concepts [58, 59, 60], relational structures [63, 64, 65, 66], and even safety-related behaviors such as refusal to generate harmful outputs [61, 62].
@@ Appendix B Extended Experimental Details and Further Results
B.1 Synthetic Experiment
The synthetic experiments were conducted using the codebase provided by Bussmann et al. [20]. Specifically, we introduced the following modifications:
  • Extended the framework to support both MP-SAE and BatchTopK variants.
  • Enforced mutual exclusivity among child features, yielding a sparsity of 1 or 2.
  • Introduced variance in the firing magnitudes of the features (see Section B.1.4).
  • Added controlled intra-level correlations between features.
Before detailing our experimental setup, we first introduce the concept of Matryoshka Sparse Autoencoders (SAEs).
@@ B.1 Synthetic Experiment
B.1.1 Dataset Generation
The tree structure contains 20 nodes, each corresponding to a unit-norm vector in $\mathbb{R}^{20}$, as illustrated in Figure 2 and detailed in Figure 10. The exact tree configuration used in our experiments is available on the project’s GitHub repository [^3] https://github.com/mpsae/MP-SAE.
Feature Activations
The root node (index 0) is fixed to the zero vector, i.e., $0\in\mathbb{R}^{20}$. All child nodes are mutually exclusive, leading to input sparsity levels of either 1 or 2. Leaf nodes—green parents without children (indices 13–20)—are independently activated with probability 0.05 ($\mathbb{P}[z_{p}>0]=0.05$). Internal green parent nodes (1, 5, and 9) are activated with probability 0.2 ($\mathbb{P}[z_{p}>0]=0.2$). Blue child nodes (2–4, 6–8, 10–12) are activated with probability 0.2 ($\mathbb{P}[z_{c}>0\mid z_{p}>0]=0.2$), conditional on their corresponding parent (1, 5, or 9) being active, in line with Definition 4.1. This results in an expected target sparsity of $\ell_{0}=1.36$.
Activation Magnitudes
If a node is active, its sparse code is drawn from a Gaussian distribution $z\sim\mathcal{N}(1.5,\nicefrac{{1}}{{4^{2}}})$. This differs from the generative process of Bussmann et al. [20], where $z_{p}$ and $z_{c}$ are fixed across all inputs, making their firing magnitudes perfectly correlated ($z_{p}=\lambda z_{c}$). We provide a justification for this design choice in Section B.1.4. The input $\bm{x}$ is then constructed as the sum of the active node vectors scaled by their corresponding code values:
$\bm{x}=z_{p}\cdot\bm{D}_{p}+z_{c}\cdot\bm{D}_{c}.$
Introducing Intra-Level Correlations.
In the original codebase, all dictionary directions are orthogonal, obtained via QR decomposition. To introduce correlations specifically between nodes at the same hierarchical level, we perturb each dictionary element as
@@ B.1.1 Dataset Generation
where the sum is restricted to elements sharing the same parent, followed by renormalization. The parameter $\varepsilon$ is tuned empirically to achieve the desired degree of intra-level correlation.
B.1.2 Training
Our training pipeline follows the same procedure as that used by Bussmann et al. [20]. Samples are drawn from the tree structure in batches of 200 over 15,000 training steps, using the Adam optimizer with $\beta=(0.5,0.9375)$ and a learning rate of $3\times 10^{-2}$. To stabilize optimization, gradient norms are clipped at 1.
For sparsity control, all SAEs, except MP-SAE, are trained to match a target $\ell_{0}$ sparsity of 1.36.
  • Vanilla and Matryoshka SAEs begin with a low $\ell_{1}$ regularization during the first 3,000 steps, which is subsequently increased or decreased based on whether the observed sparsity is below or above the target.
@@ B.1.2 Training
  • MP-SAE, in contrast, does not need to rely on a fixed sparsity target. (make it sound more like an advantage)Instead, it unrolls the encoder until either the residual norm drops below 0.05 or the support of the sparse code $z$ stabilizes—avoiding infinite loops when the residual cannot be sufficiently reduced. Unlike the other SAEs, MP-SAE uses tied weights.
B.1.3 Detailed Results on Synthetic Data
@@
All models are evaluated using a fixed sparse code $\bm{z}$, shown in Figure 11(b), which depicts the ground truth activation matrix $\bm{z}\in\mathbb{R}^{N\times p}$. When varying the degree of intra-level correlation, Vanilla SAE and BatchTopK consistently exhibit feature absorption. Matryoshka SAE sometimes alleviates this issue but does so by suppressing intra-level correlations entirely. In contrast, MP-SAE reliably recovers both the hierarchical structure and intra-level correlations at low to moderate levels (0, 0.3) and at high correlation levels (0.9), it prioritizes preserving the hierarchy over matching correlation. We summarize our detailed observations for each correlation level below:
@@
  • Figure 15 Correlation = 0.9): At high levels of intra-level correlation, all SAEs exhibit degraded performance in recovering features. Vanilla SAE and BatchTopK still capture some similarity among the last-layer children but tend to flatten the hierarchy by focusing on correlated low-level features while neglecting higher-level ones. MP-SAE, interestingly, separates each hierarchical level into two, resulting in four orthogonal sublevels composed of 1, 10, 3, and 6 features, respectively. Within each ground-truth level of correlated features, MP-SAE learns one anchor direction capturing the shared structure, and additional orthogonal directions that describe deviations from this anchor to reconstruct the remaining features, as observed in the correlation matrices $\bm{D}^{\top}\bm{D}$. Since it is not trained with a fixed target sparsity, this separation is achieved by increasing the support: most inputs now activate 2–4 components compared to the ground-truth sparsity of 1–2. This behavior preserves conditional structure even under strong entanglement—at the cost of reduced correlation—illustrating MP-SAE’s bias toward maintaining hierarchical organization when representations are highly coupled.
@@
B.1.4 Introducing Variance in Firings
In the original codebase of Bussmann et al. [20], activation magnitudes are fixed across all inputs. This implicitly assumes that parent and child firings are perfectly correlated:
$z_{p}=\lambda z_{c}.$
Such a configuration defines a strict hierarchical setting that inherently favors feature absorption. In contrast, our formulation does not impose any constraint on the correlation between firing magnitudes across concepts. Our setup generalizes the original formulation by removing the perfect correlation assumption. This is achieved by introducing variance in the activation magnitudes while maintaining the sole structural assumption that a child feature can only be active if its parent is active (Definition 4.1).
Rationale.
To illustrate the consequences of assuming perfect correlation in activation magnitudes, consider two concepts: “red object” (parent) and “apple” (child). The child can only activate when the parent does (assuming all apples are red). If the parent’s firing magnitude $z_{p}$ represents color intensity and the child’s $z_{c}$ represents size, assuming $z_{p}$ and $z_{c}$ are perfectly correlated is equivalent to stating that larger apples are always redder. This perfectly merges the two informational components—color intensity ($z_{p}\bm{D}_{p}$) and apple size ($z_{c}\bm{D}_{c}$)—into a single absorbed direction $z_{p}\bm{D}_{p}+z_{c}\bm{D}_{c}$, thereby eliminating their hierarchical distinction. In this extreme case, it becomes unclear whether the absorbed feature should still be interpreted as a child, since its activation is perfectly predictable from the parent. We argue that imperfect correlation between $z_{p}$ and $z_{c}$ is a necessary condition for preserving hierarchical interpretability.
Implications for Sparse Methods.
Following Occam’s razor, an $\ell_{0}$-penalized method such as MP-SAE will optimally encode both parent and child with a single feature when their activations are nearly identical, since the absorbed feature forms a highly compressible direction [143]. Encoding the absorbed feature $z_{p}\bm{D}_{p}+z_{c}\bm{D}_{c}$ yields a sparsity of 1, whereas disentangling parent and child requires a sparsity of 2. Therefore, it is optimal for MP-SAE to represent the parent and child through a single absorbed feature when it can perfectly reconstruct $\bm{x}$ with lower sparsity, which is precisely what occurs when no firing variance is introduced, as in the original benchmark. However, because of the greedy nature of Matching Pursuit, MP-SAE may still learn to disentangle the parent and child, depending on initialization. Figure 16 illustrates this phenomenon: it shows the correlation matrices and ground-truth alignment for the unmodified Matryoshka SAEs benchmark, trained for 15,000 iterations with a threshold of $T=0.6$ for MP-SAE.
Additional results with different child firing distributions.
In our generative data process, both parent and child activations initially follow the same distribution, $z_{p}\sim\mathcal{N}(1.5,\nicefrac{{1}}{{4^{2}}})$ and $z_{c}\sim\mathcal{N}(1.5,\nicefrac{{1}}{{4^{2}}})$. However, high variance combined with equal means ($\mu_{p}=\mu_{c}$) can produce edge cases where a strong child and a weak parent yield $\bm{x}=z_{p}\bm{D}_{p}+z_{c}\bm{D}_{c}\approx z_{c}\bm{D}_{c}$, effectively making the parent contribution negligible and thus deviating from Definition 4.1.
To assess the robustness of MP-SAE under varying child activation statistics, we evaluated its performance across a grid of child means and variances, while keeping the parent distribution fixed as $z_{p}\sim\mathcal{N}(1,\nicefrac{{1}}{{4^{2}}})$. For each configuration, we computed an absorption score, defined as the average cosine similarity between the learned child features and their corresponding ground-truth parents over 15 independent runs. Lower similarity indicates successful disentanglement, whereas higher similarity reflects feature absorption. NaN values correspond to cases where the model failed to recover the child feature due to its very low firing amplitude. All activations were constrained to remain positive when active (entries marked “–” in Table 1 denote untested configurations).
Across all evaluated conditions, MP-SAE consistently achieves low absorption scores, confirming its robustness and its ability to recover hierarchical structure even under challenging firing regimes. Indeed, when the child average firing magnitude is particularly low (0.1), MP-SAE remains the only method capable of recovering the hierarchical structure, owing to its iterative greedy encoder. Interestingly, when the variance of the child distribution decreases for a mean of 0.50.5—a regime where overall performance becomes comparable to Matryoshka—the two methods exhibit opposite trends: Matryoshka performs better at low variance, whereas MP-SAE improves as variance increases. This observation suggests that the two approaches embody complementary inductive biases: Matryoshka is particularly effective in scenarios dominated by strong feature absorption and near-perfect correlation, while MP-SAE excels when moderate variance is present in the firing magnitudes.
Figure 16: Evaluating MP-SAE on the original Matryoshka benchmark [^5] https://github.com/noanabeshima/matryoshka-saes, where firings are perfectly correlated. For each SAE, four independent runs are shown. The top row displays $\bm{D}^{\top}\bm{D}$ (self-similarity of learned features), and the bottom row shows $\bm{D}_{\text{GT}}^{\top}\bm{D}$ (alignment with ground truth). Top left: ground-truth correlation matrix, where features are perfectly orthogonal. Matryoshka overcomes feature absorption at the cost of negative interference. Vanilla consistently learns absorbed features. MP-SAE occasionally recovers parent–child pairs thanks to its inductive bias, but under perfect correlation, its behavior is unstable.
Table 1: Comparison of absorption scores across sparse autoencoders for different child activation distributions. Child activations follow normal distributions with varying means (rows) and standard deviations (columns), while the parent distribution is fixed as $z_{p}\sim\mathcal{N}(1,\nicefrac{{1}}{{4^{2}}})$. Lower absorption scores indicate better disentanglement between parent and child features. MP-SAE maintains low absorption across settings and remains the only method robust to low mean child firings.
MP 0.1 0.5 1.0
1/400 3.6e-03 7.3e-02 1.3e-03
1/40 5.0e-03 6.2e-02 1.2e-03
1/8 4.3e-02 1.1e-03
1/4 1.1e-03
Matryoshka 0.1 0.5 1.0
1/400 nan 4.8e-02 2.1e-01
1/40 nan 9.0e-02 1.5e-01
1/8 9.6e-02 2.1e-01
1/4 2.2e-01
Vanilla 0.1 0.5 1.0
1/400 nan 7.0e-01 4.5e-01
1/40 nan 7.0e-01 4.5e-01
1/8 7.0e-01 4.7e-01
1/4 4.7e-01
BatchTopK 0.1 0.5 1.0
1/400 nan 5.0e-01 4.8e-01
1/40 nan 5.4e-01 4.5e-01
1/8 4.9e-01 4.6e-01
1/4 4.2e-01